OpenStudy (istokes760):

Find the difference. (16x3 + 2x2) − 4x3 =

OpenStudy (phi):

you have 16 x^3 's and take away 4 x^3 's how many x^3 's do you have left ?

OpenStudy (amorfide):

by the commutative rule, addition and subtraction can go in any order, so a+b+c = a+(b+c) = (a+b)+c here we have (16x3 + 2x2) − 4x3 = (16x3 - 4x3) + 2x2 so just add them up for example if I had 3a3 - 2a3 this would give a3

OpenStudy (istokes760):

i suck at algebra so soory if i dont get it at first. okay so yall saying to just add them. this is what i got 16x3+2x2=18x2-4x2 and this equals 12x2

OpenStudy (phi):

almost. you can combine (add or subtract) the same things like 2 cows plus 1 cow is 3 cows but 2 cows plus 1 horse is not 3 cows (nor 3 horses), it has to stay 2 cows + 1 horse

OpenStudy (phi):

does that idea make sense ? if so, you next need to know "what are the same thing" in algebra if you have the same letter "x" as in 2x + 1x (or just 2x + x) that is the same thing a different letter would be different: x+y can't be simplifed the other way to be different: x*x (x times x) is *not* the same as x or in general x*x*x is not the same as x or x*x we usually use the "little number in the upper right" to write x*x*x as \( x^3\)

OpenStudy (phi):

in your problem \[ (16x^3 + 2x^2) − 4x^3 = \] you have \[ 16x^3 - 4x^3 \] can you simplify that ?

OpenStudy (istokes760):

yes i can thank you so much for the help!!!

OpenStudy (phi):

you should bet 12x^3 and you also have + 2x^2 but x^3 's are different from x^2 so you are done: \[ 12x^3 + 2x^2 \]

OpenStudy (phi):

*should get

OpenStudy (istokes760):

thanks this was a big help since im taking online classes